Jozy Altidore invested $8,600 at 9% annual interest, and left the money invested without withdrawing any of the interest for 13 years. At the end of the 13 years, Jozy withdrew the accumulated amount of money.

Click here to view the factor table.

(For calculation purposes, use 5 decimal places as displayed in the factor table provided.)

(a) What amount did Jozy withdraw, assuming the investment earns simple interest?

The amount Jozy withdrew

$enter the accumulated amount assuming the investment earns simple interest


(b) What amount did Jozy withdraw, assuming the investment earns interest compounded annually? (Round answer to 2 decimal places, e.g. 25.25.)

The amount Jozy withdrew

$enter the accumulated amount rounded to 2 decimal places assuming the investment earns interest compounded annually
Answer #1

a) Amount of withdraw with Simple interest = 8600+(8600*9%*13) = 18662

b) Amount of withdraw with compound interest = 8600*3.06580 = $26365.88